40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1334     Entry time: Tue Feb 24 02:23:40 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Locking - MC board bad 
Rob, Yoichi, Alberto, Kiwamu, Kakeru

We found that the OMC alignment feedback was on for the POS X loop even though the OMC was not locked.
This caused the PZT mirror to be tilted in yaw a lot. This was probably the reason for the mysterious shift in the AS beam last week, because the AS RF beam is picked up after this PZT mirror.
Rob aligned the OMC and we re-centered the AS PDs and the CCD.
This changed the DARM RF gain, so we changed it from 3 to 1. This gain used to be -1. It is still not understood why the polarity was changed.
The MC length was changed ? We should check the sideband transmission.

After this, we reached to the arm power 4. But the IFO loses lock immediately after the moving zero is turned off.
At this stage, the CARM loop bandwidth is supposed to be high enough that the moving zero is no longer necessary.
However, when we measured the MCL loop gain with several different AO path gains, the loop shape did not change at all.
This led us to suspect the AO path may not be connected. The cabling from the common mode board to the MC board seemed ok.
We tested the signal flow in the MC board using a signal generator and an oscilloscope.
Then we found that a signal injected to the IN2 (AO path) does not reach to the TP1A (right after the boost stages), though the signal is visible in the OUT2 (monitor BNC right after the initial amplifier (B-amp) for the AO path). The signal from IN1 (MC REFL) can be observed at TP1A. This means something is broken between the B-amp and the sum-amp in the AO path.
We will check the MC board tomorrow.
ELOG V3.1.3-